OpenStudy (anonymous):

@myininaya I know what I need to do and all but I just need you to tell me if I'm on the right path and if not then I will just look over on the word problem and see what I did wrong. A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent.

OpenStudy (anonymous):

@myininaya if you can, please come

myininaya (myininaya):

\[g(x)=200(1.03)^{2x}?\]

OpenStudy (anonymous):

ok so I have an idea on how im supposed to do this so I want you to tell me if im on the right path. Ok?

myininaya (myininaya):

do you know law of exponents ?

myininaya (myininaya):

alright

myininaya (myininaya):

shoot

OpenStudy (anonymous):

So what I think I need to do is go and multiply 1.03 and 2, then with 200

OpenStudy (anonymous):

is it ok? if not i will go over.

myininaya (myininaya):

well Kay let me remind of one particular law we will need for this

myininaya (myininaya):

\[(a^{xy})=(a^y)^x \]

myininaya (myininaya):

pretend y is 2 if you want

OpenStudy (anonymous):

? so i have to multiply 1.03 by 2 and that's it or is it more complicated than that.

myininaya (myininaya):

no something like a^2 means a*a

myininaya (myininaya):

not 2*a

myininaya (myininaya):

\[g(x)=200(1.03)^{2x} \\ g(x)=200(1.03^2)^x\]

myininaya (myininaya):

1.03^2=?

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

yeah i forgot that it is an exponent not just a number so it would be 1.03 * 1.03

myininaya (myininaya):

yes for the base yes

OpenStudy (anonymous):

then by 200 and my answer would be there

OpenStudy (anonymous):

right?

myininaya (myininaya):

well 200 is on the outside

myininaya (myininaya):

you just need to do this: \[g(x)=200(1.03)^{2x} \\ g(x)=200(1.03^2)^x \\ g(x)=200(1.03 \cdot 1.03)^x\]

myininaya (myininaya):

unless g was meant to be the following: \[g(x)=(200 \cdot 1.03)^{2x}\]?

OpenStudy (anonymous):

oh yeah cuz I did not know if i wasd supposed to leave it like that or multiply by 200 afterwards. That's why i wanted someone to go over my original thoughts

myininaya (myininaya):

yep the only step you need to do is do the 1.03*1.03

myininaya (myininaya):

you cannot do anything with the 200 because it is outside the ( )^x part

myininaya (myininaya):

so you will just bring it down

myininaya (myininaya):

\[g(x)=200(?)^x\] what did you get for the question mark

OpenStudy (anonymous):

well thanks for checking.

OpenStudy (anonymous):

ill tell you... 1.0609

myininaya (myininaya):

that's good once you replace that ? with that number you are done

myininaya (myininaya):

don't do anything else

myininaya (myininaya):

\[g(x)=200(1.0609)^x \]

OpenStudy (anonymous):

thank you, i gave you a medal

myininaya (myininaya):

I gave you a medal for your effort and wanting to learn how to do the problem. You are awesome @KayleeNeedsHelp ! :)

OpenStudy (anonymous):

well thanks! I am so glad your there to help!

OpenStudy (anonymous):

is there a way i can call you for short because its harder with your user. Thanks again, bye!

myininaya (myininaya):

lol

myininaya (myininaya):

you can private message me a link to your problem if that makes it easier

myininaya (myininaya):

in the future if you want

OpenStudy (anonymous):

oh yeah sure i'll message you in any case

myininaya (myininaya):

but unfortunately I don't think there are any shortcuts to call people

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

See ya later

myininaya (myininaya):

see ya